### v2

# Learning to Optimize for Mixed-Integer Non-linear Programming

Abstract
Mixed-integer non-linear programs (MINLPs) arise in various domains, such as energy systems and transportation, but are notoriously difficult to solve. Recent advances in machine learning have led to remarkable successes in optimization tasks, an area broadly known as learning to optimize. This approach includes using predictive models to generate solutions for optimization problems with continuous decision variables, thereby avoiding the need for computationally expensive optimization algorithms. However, applying learning to MINLPs remains challenging primarily due to the presence of integer decision variables, which complicate gradient-based learning. To address this limitation, we propose two differentiable correction layers that generate integer outputs while preserving gradient information. Combined with a soft penalty for constraint violation, our framework can tackle both the integrality and non-linear constraints in a MINLP. Experiments on three problem classes with convex/non-convex objective/constraints and integer/mixed-integer variables show that the proposed learning-based approach consistently produces high-quality solutions for parametric MINLPs extremely quickly. As problem size increases, traditional exact solvers and heuristic methods struggle to find feasible solutions, whereas our approach continues to deliver reliable results. Our work extends the scope of learning-to-optimize to MINLP, paving the way for integrating integer constraints into deep learning models. Our code is available at https://github.com/pnnl/L2O-pMINLP.
1 Introduction
Mixed-integer optimization is fundamental to a broad spectrum of real-world applications spanning problems in fields as diverse as pricing Kleinert et al. (2021), battery dispatch (Nazir & Almassalkhi, 2021), transportation (Schouwenaars et al., 2001), and optimal control (Marcucci & Tedrake, 2021). These problems involve discrete decisions, such as determining the number of items or the activation of generators, combined with complex non-linear system constraints. Mixed-integer linear programming (MILP) has been widely adopted due to its well-established solution techniques. However, many practical problems exhibit non-linear relationships, leading to mixed-integer non-linear programs (MINLPs). Unlike MILPs, where techniques such as branch-and-bound (Land & Doig, 2010), cutting planes (Gomory, 2010), and heuristics (Crama et al., 2005; Johnson & McGeoch, 1997) have matured, MINLPs require more complex approaches due to the combination of discrete variables and non-convex constraints and objective function. Standard methods include outer approximation (Fletcher & Leyffer, 1994), spatial branch-and-bound (Belotti et al., 2009), and decomposition techniques (Nowak, 2005), but these often struggle to scale to large problems.
Many applications demand that MINLPs be solved within a limited time budget, further complicating the picture. To overcome this, learning-to-optimize (L2O) methods offer a promising alternative by leveraging machine learning (ML) to enhance or even replace conventional optimization approaches. In particular, end-to-end optimization directly maps input instance parameters to solutions of optimization problems through a trained model (Kotary et al., 2021a; Chen et al., 2022). By identifying patterns in a distribution of similar instances of the same optimization problem and predicting solutions accordingly, end-to-end optimization can bypass traditional, computationally intensive optimization methods, enabling faster computation and improved scalability.
Many real-world applications have stringent requirements on operational, physical, or safety constraints. Thus, recent research in machine learning has focused on the feasibility issue. While various strategies exist, such as embedding hard constraints into neural network architectures (Hendriks et al., 2020), using penalty terms in loss functions for soft constraints (Pathak et al., 2015; Jia et al., 2017), or projecting solutions onto feasible regions (Donti et al., 2021), these methods are not directly applicable for problems that involve integer decisions.
This work tackles, for the first time, the non-differentiability associated with predicting integer variables using a deep neural network, in conjunction with non-linear objective and constraints. This challenge has been underexplored in learning-based methods due to the absence of useful gradient information. To that end, we propose two differentiable correction layers for rounding, allowing for gradient-based optimization of a neural network that generates high-quality integer solutions while maintaining feasibility. Our contributions are as follows:
  • We study the learning-to-optimize problem in the context of parametric MINLP, enabling, for the first time, quick, dynamic solution generation as problem parameters change.
  • We propose two novel differentiable correction layers that effectively handle the non-differentiability of integer outputs in deep learning models. Combined with a penalty method for soft constraint satisfaction, we are able to learn a neural network mapping from instance parameters to solutions through gradient-based learning. Not only does this yield an extremely fast heuristic at test time, but the method is also self-supervised and thus efficiently trainable. Additionally, we introduce a strategy of progressively increasing penalties to further improve feasibility in the generated solutions.
  • We conduct extensive experiments on three problem classes from the literature: a convex integer quadratic problem, a non-convex integer problem, and a non-convex mixed-integer problem. Our learning-based methods consistently yield high-quality, feasible solutions extremely fast, outperforming exact solvers and heuristics. For the largest test instances we consider, the baselines fail to produce any solutions, whereas our methods continue to generate good, feasible solutions in most instances.
2 Related Work
End-to-end optimization.
End-to-end optimization focuses on training machine learning models to predict the problem solutions, bypassing the need for computationally expensive solvers. One of the early approaches was proposed by Hopfield & Tank (1985), who used Hopfield networks to solve the traveling salesperson problem by incorporating a Lagrangian penalty for constraint feasibility. Similarly, Fioretto et al. (2020) applied the Lagrangian penalty in the context of continuous non-linear optimization for energy systems. In addition to penalty-based methods for ensuring feasibility, Pan et al. (2020) embedded certain constraints directly into neural networks by leveraging the range of output values and solving linear systems. Although these supervised learning methods significantly reduce inference time, they typically require large offline datasets of solutions (Gleixner et al., 2021; Kotary et al., 2021b), which can be impractical for large-scale problems where generating solutions is computationally expensive. This limitation highlights the need for self-supervised learning approaches (Donti et al., 2021), which minimize both the objective function and constraint violation from the predicted values, without relying on the imimitation of pre-solved solutions. Our method first extends this self-supervised paradigm to problems involving discrete decision variables, further broadening its applicability to mixed-integer optimization.
Constrained neural architectures.
Specific neural network architectures can be designed to impose certain classes of hard constraints. For instance, Hendriks et al. (2020) incorporate linear operator constraints directly into the model design. Vinyals et al. (2015) and Dai et al. (2017) leveraged the inherent structure of graphs to construct feasible solutions for the traveling salesperson problem. Additionally, Kervadec et al. (2022) demonstrated that employing a log-barrier method for inequality constraints improves accuracy, constraint satisfaction, and training stability. Penalty methods (Pathak et al., 2015; Jia et al., 2017), which impose inequality constraints through regularization terms in the loss function, have also gained popularity for constraining neural networks. As noted by Márquez-Neila et al. (2017), in practice, methods that incorporate hard constraints rarely outperform their soft constraint counterparts, despite the latter offering weaker theoretical performance guarantees. Building on penalty methods, Donti et al. (2021) proposed a differentiable correction approach to complete partial solutions for linear equations and project solutions onto the feasible region. In this paper, we adopt a penalty method for handling constraints and introduce two novel differentiable rounding correction layers to guarantee the integrity of the solution.
Learning for mixed-integer programming.
There has been significant interest in using ML to accelerate the solution of integer programs. The vast majority of the work in this space focuses on learning search strategies for exact MILP solvers. This includes parameter tuning (Xu et al., 2011), preprocessing (Berthold & Hendel, 2021), branching variable selection (Khalil et al., 2016; Alvarez et al., 2017; Gasse et al., 2019; Zarpellon et al., 2021), node selection (He et al., 2014), heuristic selection (Chmiela et al., 2021), and cut selection and generation (Deza & Khalil, 2023). Another line of research in ML-for-MILP relates to learning to generate integer solutions heuristically (Nair et al., 2020; Khalil et al., 2022; Ding et al., 2020; Sonnerat et al., 2021; Song et al., 2020; Bertsimas & Stellato, 2022; Huang et al., 2023; Ye et al., 2024). We refer to the surveys of Bengio et al. (2021) and Zhang et al. (2023) for more details. In contrast, there has been much less work on MINLP. Illustrative examples include the work of Cauligi et al. (2022) who proposed a two-stage algorithm for quickly finding high-quality solutions for mixed-integer convex programs (MICPs), Baltean-Lugojan et al. (2019) who use supervised learning to select cuts for quadratic optimization, Nowak et al. (2018) who learn to solve quadratic assignment problems with graph networks, and Bonami et al. (2022) who use a classifier to decide on the linearization of mixed-integer quadratic problems. Most relevant to our method is the recently proposed SurCO approach of Ferber et al. (2023). They focus on mixed-integer problems with non-linear objective and linear constraints, learning to approximate the former with a linear function for a simpler heuristic optimization. Our approach differs from all of the above in its scope, addressing the most general class of MINLPs.
Differentiable optimization.
A different category of methods integrates optimization solvers as layers within deep neural network architectures (Agrawal et al., 2019). These methods can handle various types of optimization problems, such as quadratic programs (Amos & Kolter, 2017; Sambharya et al., 2023), stochastic optimization (Donti et al., 2017), submodular optimization (Djolonga & Krause, 2017), and even integer linear programs (Wilder et al., 2019; Berthet et al., 2020; Pogančić et al., 2020). In these approaches, optimization algorithms or solvers are embedded within the neural network, allowing gradients of optimization solvers to be computed and propagated during backpropagation. King et al. (2024) shows how differentiable optimization can enhance the convergence of proximal operator algorithms via end-to-end learning of proximal metrics. However, as Tang & Khalil (2024) noted, training with a differentiable optimizer requires iteratively solving optimization throughout the training process, making the computational burden prohibitively expensive. In contrast, our self-supervised approach generates solutions directly through neural network structures, eliminating the need to repeatedly call high-complexity solvers and thus significantly reducing computational overhead.
3 Learning to Optimize MINLPs: a Problem Formulation
A generic learning-to-optimize formulation for parametric mixed-integer non-linear programming is given by:
$\min_{\Theta}\quad$ $\mathbb{E}_{{\bm{\xi}}\sim\mathcal{P}_{\xi}}\,{\bf f}({\bf x^{\bm{\xi}}},{\bm{\xi}})$ (1a)
s.t. ${\bf g}({\bf x^{\bm{\xi}}},{\bm{\xi}})\leq 0,$ $\forall{{\bm{\xi}}\in\mathcal{P}_{\xi}}$ (1b)
${\bf x^{\bm{\xi}}}\in\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}},$ $\forall{{\bm{\xi}}\in\mathcal{P}_{\xi}}$ (1c)
${\bf x^{\bm{\xi}}}=\bm{\psi}_{\Theta}({\bm{\xi}}),$ $\forall{{\bm{\xi}}\in\mathcal{P}_{\xi}}$ (1d)
Here, $\mathcal{P}_{\xi}$ represents the distribution over parametric MINLP instances and ${\bm{\xi}}\in\mathbb{R}^{n_{\xi}}$ is a vector of instance parameters. Vector ${\bf x^{\bm{\xi}}}\in\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ represents the mixed-integer decision variables for parameters ${\bm{\xi}}$. The mapping $\bm{\psi}_{\Theta}({\bm{\xi}})$ is a neural network with weights $\Theta$ that outputs a parametric solution ${\bf x^{\bm{\xi}}}$ for parameters ${\bm{\xi}}$. The goal is to find the neural network weights that minimize the expected objective function value in eq. 1a over the parameter distribution, subject to the constraints in eq. 1b. Note that $\bf{g}(\cdot)$ is a vector-valued function representing one or more inequality constraints. As is typical in MINLP, we assume that the objective and constraint functions are differentiable.
As is typical, we will train the neural network using empirical risk minimization on a sample of $m$ training instances drawn from the parameter distribution $\mathcal{P}_{\xi}$. Then, the average value of the objective function ${\bf f}(\cdot)$ serves as a natural loss function. Our approach is self-supervised since the loss calculation does not require any labeled data. This is particularly appealing as computing optimal or even feasible solutions to a MINLP is in general extremely challenging. Solely minimizing the average objective is insufficient if the solutions violate the constraints. Therefore, similarly to Donti et al. (2021), we incorporate penalty terms into the loss function to account for constraint violations, enhancing the feasibility of the solution and resulting in a soft-constrained empirical risk minimization loss function given:
$\mathcal{L}(\Theta)=\frac{1}{m}\sum_{i=1}^{m}\Big{(}{\bf f}({\bf x}^{i},{\bm{\xi}}^{i})+\lambda\cdot\|\left({\bf g}({\bf x}^{i},{\bm{\xi}}^{i})\right)_{+}\|_{2}^{2}\Big{)}\text{ with }{\bf x}^{i}=\bm{\psi}_{\Theta}({\bm{\xi}}^{i})\;\quad\forall i\in[m],$ (2)
where $m$ is the number of sampled parametric scenarios in the training dataset, $\left(\cdot\right)_{+}$ ensures only positive constraint violations are penalized (implemented via a ReLU function), $\|\cdot\|_{2}$ denotes the Euclidean norm, and $\lambda>0$ is a hyperparameter that balances the trade-off between minimizing the objective function and satisfying the constraints.
4 Preliminaries: Differentiating through Discrete Operations
Straight-through Estimator.
The Straight-through Estimator (STE) (Bengio et al., 2013) is a simple yet effective method for handling non-differentiable operations in neural networks. In our approaches, STE plays a crucial role in enabling backpropagation through discrete operations. During the forward pass, STE applies a (non-differentiable) discrete operation, such as rounding a variable up or down, binarizing it, or using an indicator function $\mathbb{I}(\cdot)$. However, in the backward pass, STE replaces the non-existent gradient of these discrete functions with soft approximations. For rounding operations, the gradient of the identity function is used during backpropagation, whereas for binarization or indicator functions, the gradient of the Sigmoid function is applied.
Gumbel-Sigmoid Noise.
Although the STE is effective for backpropagating through discrete decisions, it lacks the stochasticity that can improve model training. This is where the Gumbel-noise method (Jang et al., 2016) comes into play. Specifically, Gumbel noise perturbs the logits before applying the Sigmoid function, allowing for randomness in the binary decisions. After this, a hard binarization step is applied using the STE, ensuring that the final outputs are discrete binary values while retaining gradients for backpropagation. Further technical details can be found in Appendix A.
5 Learning to Optimize MINLPs with Correction Layers
Figure 1: Conceptual diagram for our method. self-supervised differentiable programming-based solution approach for parametric MINLP problems.
Algorithm 1 Learning-to-optimize MINLPs with Correction Layers: Forward Pass.
1:Training instance ${{\bm{\xi}}^{i}}$, neural networks $\pi_{\Theta_{1}}(\cdot)$ and $\delta_{\Theta_{2}}(\cdot)$
2:Predict a continuously relaxed solution $\bar{\bf x}^{i}\leftarrow\pi_{\Theta_{1}}({{\bm{\xi}}^{i}})$
3:Obtain an initial correction prediction ${\bf h}^{i}\leftarrow\delta_{\Theta_{2}}(\bar{\bf x}^{i},{\bm{\xi}}^{i})$
4:Update continuous variables: $\hat{{\bf x}}^{i}_{r}\leftarrow\bar{{\bf x}}^{i}_{r}+{\bf h}^{i}_{r}$
5:Round integer variables down: $\hat{{\bf x}}^{i}_{z}\leftarrow\lfloor\bar{{\bf x}}^{i}_{z}\rfloor$
6:if using Rounding Classification then
[…]
8:else if using Learnable Threshold then
9:     Compute ${\bf v}^{i}\in[0,1]^{n_{z}}\leftarrow\text{Sigmoid}({\bf h}_{z}^{i})$
10:     Compute rounding direction: ${\bf b}^{i}\leftarrow\mathbb{I}\big{(}(\bar{{\bf x}}^{i}_{z}-\hat{{\bf x}}^{i}_{z})-{\bf v}^{i}>0\big{)}$
11:end if
12:Update integer variables: $\hat{{\bf x}}^{i}_{z}\leftarrow\hat{{\bf x}}^{i}_{z}+{\bf b}^{i}$
13:return $\hat{\bf x}^{i}$
We are now ready to introduce the two differentiable correction layers designed to handle the integrality constraints of MINLPs. They differ in how they round an integer variable’s fractional value but are equally easy to train with gradient descent and fast at test time. We decompose the mapping $\psi_{\Theta}:\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ from an instance parameter vector to a candidate mixed-integer solution into two steps.
  1. 1.
    The first step consists in applying a learnable relaxed solution mapping $\pi_{\Theta_{1}}:\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}+n_{z}}$ encoded by a deep neural network with weights $\Theta_{1}$. It outputs a continuously relaxed solution $\bar{\bf x}^{i}\in\mathbb{R}^{n_{r}+n_{z}}$ without enforcing the integrality requirement. Note that continuous variables are also predicted in this first step.
  2. 2.
    The second step is a differentiable correction layer $\varphi_{\Theta_{2}}:\mathbb{R}^{n_{r}+n_{z}}\times\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}}\times\mathbb{Z}^{n_{z}}$ that takes as input the instance parameter vector and the continuous solution produced in the first step, and outputs a candidate mixed-integer solution while maintaining differentiability. Here, $\Theta_{2}$ represents the weights of the neural network $\delta_{\Theta_{2}}:\mathbb{R}^{n_{r}+n_{z}}\times\mathbb{R}^{n_{\xi}}\mapsto\mathbb{R}^{n_{r}+n_{z}}$, which is applied within the correction layer $\varphi_{\Theta_{2}}$.
Finally, during training, the loss function eq. 2 is applied to the neural network weights $\Theta=\Theta_{1}\cup\Theta_{2}$, implicitly taking into account the objective function value and constraint violations of the predicted mixed-integer solution $\hat{\bf x}^{i}$. This process is illustrated in Figure 1.
The following subsections describe two distinct approaches for designing the correction layer $\varphi_{\Theta_{2}}$; the same network $\pi_{\Theta_{1}}$ is used in both approaches. Algorithm 1 summarizes both of our approaches. Line 1 invokes the first step’s network $\pi_{\Theta_{1}}$ and lines 2–11 describe both versions of $\varphi_{\Theta_{2}}$. Our correction layers are simple and efficiently computable. While the STE and Gumbel-Sigmoid techniques have been used to train binarized or quantized neural networks, they have not been leveraged in the context of learning-to-optimize to our knowledge. As we will see in the experimental results, the simplicity of the correction layers is key to fast solution generation in large-scale MINLP problems.
These approaches can be viewed as an end-to-end learnable version of the Relaxation Enforced Neighborhood Search (RENS) algorithm (Berthold, 2014). Instead of explicitly searching the neighborhood of the relaxed solution, the neural network implicitly learns the corrections required to achieve a feasible integer solution by exploring the solution space near the integer variables while updating the continuous variables.
Rounding Classification.
Line 6 of Algorithm 1 is the key step in the rounding classification (RC) approach. For the integer variables of a problem instance, denoted with the subscript $z$, RC applies a stochastic soft-rounding to the output ${\bf h}^{i}_{z}$ of the neural network $\delta_{\Theta_{2}}(\bar{\bf x}^{i},{\bm{\xi}}^{i})$, yielding ${\bf b}^{i}\in\{0,1\}^{n_{z}}$. An entry of ${\bf b}^{i}$ determines whether the continuously relaxed value $\bar{\bf x}^{i}_{z}$ of the corresponding variable is rounded down or up. In the backward pass, STE is used in line 4 for the rounding down operation. In line 6, the derivative of the Sigmoid function is used.
Learnable Threshold.
The key steps of the learnable threshold (LT) approach are described in lines 8 and 9 of Algorithm 1. Rather than use Gumbel-Sigmoid for the rounding as in RC, LT learns a vector of per-variable rounding thresholds, ${\bf v}^{i}\in[0,1]^{n_{z}}$, that is generated by the Sigmoid in line 8 of Algorithm 1. A variable is rounded up if the fractional part of its relaxed value, $(\bar{{\bf x}}^{i}_{z}-\hat{{\bf x}}^{i}_{z})$, exceeds the threshold. The indicator function $\mathbb{I}(\cdot)$ in line 9 produces a binary output in the forward pass. In the backward pass, the gradient is approximated by that of the Sigmoid function with a slope:
${\bf b}^{i}\leftarrow\frac{1}{1+\exp\big{(}-10\cdot(\bar{{\bf x}}^{i}_{z}-\hat{{\bf x}}^{i}_{z}-{\bf v}^{i})\big{)}}.$
Here, the slope is set to 10 to sharpen the Sigmoid function.
Illustration.
Figure 2: The relaxed solutions $\bar{x},\bar{y}$ and the rounding solutions $\hat{x},\hat{y}$ across different epochs of training for the same sample instance using the Rounding Classification approach.
Figure 2 shows the evolution of both the relaxed and rounded solutions, $(\bar{x},\bar{y})$ and $(\hat{x},\hat{y})$, across different epochs of the training of an RC model on two-dimensional mixed-integer Rosenbrock problems defined as follows:
$\min_{x\in\mathbb{R},y\in\mathbb{Z}}$ $\quad(a-x)^{2}+50(y-x^{2})^{2}$
subject to $\quad y\geq{b}/{2},\quad x^{2}\leq b,\quad x\leq 0,\quad y\geq 0.$
In this formulation, $x$ is a continuous decision variable and $y$ is an integer decision variable, subject to linear constraints. The instances have parameters $a$ and $b$, which represent the input features to the neural network; for the instance illustrated in Figure 2, these are set to $3.83$ and $6.04$, respectively.
The illustration shows that the training of this differentiable rounding approach converges remarkably well in this particular instance, with the final rounding solution being very close to the optimum. We will show this to be a generalizable phenomenon, with both of our learning approaches converging to highly accurate neural network models on a variety of problem classes and sizes.
Ablation Study.
The two baselines involve only some parts of our full pipeline and thus serve as ablations. Rounding after Learning (RL) is used as a baseline, applying rounding to the relaxed solution generated by learning-based methods rather than learning-to-round as our correction layers do. In Rounding with STE (RS), the continuous values predicted by the first neural network $\pi_{\Theta_{1}}$ are directly rounded without any learned correction mechanism, i.e., without using the correction network $\varphi_{\Theta_{2}}$; STE is used to backpropagate through the rounding operator. For further details, refer to Appendix C.
6 Experimental Results
Methods.
Table 1 provides an overview of all the methods used in the following experiments. A 60-second time limit is enforced for all methods and problems. The experiments evaluate our learning-based methods, Rounding Classification (RC) and Learnable Threshold (LT), against traditional exact optimization (EX), which can compute optimal solutions but is often computationally expensive, and heuristic-based approaches such as Rounding after Relaxation (RR) and root node solutions (N1), which offer faster results without quality guarantees. Note that baselines EX and N1 include a wide range of heuristics that are embedded in the MINLP solver of choice (Gurobi or SCIP) and that are executed in conjunction with the tree search procedure; we are also implicitly comparing to these heuristics, not just to the exact search. As such, the competing methods cover a broad spectrum of optimization strategies, from exact solvers to fast heuristics, allowing for a comprehensive evaluation of solution quality and computational efficiency. In addition, two ablation baselines are included: Rounding after Learning (RL), which applies direct rounding to the relaxed solution generated by the learning-based method without any corrective adjustments, and Rounding with STE (RS), which uses the Straight-Through Estimator to handle non-differentiable rounding but without a learning-based correction layer.
Table 1: Summary of Methods. Methods with “*” use a trained model.
Method Abbr Description
Rounding Classification* RC* Learning-based rounding approach using classification for integer variable rounding.
Learnable Threshold* LT* Learning-based method where a neural network learns the threshold for rounding integer variables.
Rounding after Learning* RL* Learning-based method that learns from the relaxed problem, then rounds the inferred solution to the nearest integer.
Rounding with STE* RS* Learning-based method where the rounding process is differentiable with STE but does not involve learning.
Exact Optimization EX Solves the problem using Gurobi for convex problems and SCIP + Ipopt for non-convex problems.
Rounding after Relaxation RR Solves the continuous relaxation, then rounds the continuous solution to the nearest integer.
Root Node Solution N1 A feasible solution obtained at the root node, which uses heuristics after continuous relaxation with cutting planes.
Problem classes.
We tested the methods on a variety of optimization problems, including integer convex quadratic problems, simple integer non-convex problems, and high-dimensional mixed-integer Rosenbrock problems. These problem classes were selected to cover both convex and non-convex scenarios and to evaluate the scalability of the methods in higher-dimensional settings. Each method was assessed in terms of objective value, constraint violation, and solving time, providing a comprehensive view of their performance across different types of problems. In addition, we performed experiments on integer linear programs in Appendix E; these experiments simply serve to show that our methods could be applied to the integer linear case, though the use of MILP solvers may be preferable.
Training protocol.
The solution mapping $\pi_{\Theta_{1}}$ used across all learning-based methods (RC, LT, and RL) and the rounding correction network $\varphi_{\Theta_{2}}$ for RC and LT are based on fully connected layers with ReLU activations. Further details regarding the network hyperparameters can be found in Appendix B. For all problems, the training samples 8,000 instances from the distribution, and the test set includes 100 instances. An additional set of 1,000 instances was used for validation to fine-tune the models and select hyperparameters.
Computational setup.
All experiments were conducted on a system with an AMD Ryzen 7 5800HS CPU, 15GB of DDR4 RAM, and an NVIDIA GeForce GTX 1650 GPU. The software environment was configured with Python 3.10.13, PyTorch 2.0.1+cu117 (Paszke et al., 2019) for deep learning models, and NeuroMANCER 1.5.1 (Drgona et al., 2023) for modeling parametric constrained optimization problems.
Gurobi 11.0.3 (Gurobi Optimization, LLC, 2021) is used as the exact method for convex quadratic problems; beyond quadratic polynomials, Gurobi needs to approximate non-linearities using piecewise-linear functions. For those more general mixed-integer non-convex problems, we use SCIP 8.1.0 (Bestuzheva et al., 2021) with Ipopt 3.14.13 (Wächter & Biegler, 2006) as the continuous non-linear solver. Note that Gurobi and SCIP are considered to be among the state-of-the-art solvers for MINLP, as noted by Lundell & Kronqvist (2022) who performed a comprehensive benchmarking of more than ten MINLP solvers: “It is clear, however, that the global solvers Antigone, BARON, Couenne and SCIP are the most efficient at finding the correct primal solution when regarding the total time limit. […] Gurobi also is very efficient when considering that it only supports a little over half of the total number of problems!”
6.1 Convex Quadratic Problem
Table 2: Result for a Convex Quadratic Problem. Each problem size is evaluated on a test set of 100 instances. “Obj Mean” and “Obj Median” represent the mean and median objective values for this minimization problem, with smaller values being better. “Infeasible” denotes the fraction of infeasible solutions, and “Time (Sec)” is the average solving/inference time per instance. The “—” symbol indicates that no solution is found for any instance within 60 sec.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500
RC Obj Mean $0.608$ $-1.602$ $-4.148$ $-13.109$ $-13.589$ $-31.206$ $-73.570$
Obj Median $0.413$ $-1.600$ $-4.194$ $-13.131$ $-13.587$ $-31.247$ $-73.761$
% Infeasible $3\%$ $2\%$ $4\%$ $1\%$ $4\%$ $5\%$ $8\%$
Time (Sec) $0.0026$ $0.0031$ $0.0038$ $0.0031$ $0.0039$ $0.0044$ $0.0074$
LT Obj Mean $0.649$ $-1.818$ $-4.329$ $-12.905$ $-13.745$ $-31.261$ $-73.457$
Obj Median $0.470$ $-2.078$ $-4.341$ $-12.946$ $-13.747$ $-31.320$ $-73.414$
% Infeasible $5\%$ $7\%$ $1\%$ $2\%$ $7\%$ $1\%$ $8\%$
Time (Sec) $0.0032$ $0.0030$ $0.0038$ $0.0031$ $0.0045$ $0.0045$ $0.0098$
EX Obj Mean $0.294$ $-2.779$ $-5.120$ $-15.885$
Obj Val Median $0.129$ $-2.991$ $-5.130$ $-15.940$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1451$ $0.1996$ $5.4154$ $61.2890$
RR Obj Mean $0.211$ $-2.858$ $-5.179$ $-16.173$ $-21.922$ $-46.727$ $-106.526$
Obj Median $0.058$ $-3.033$ $-5.217$ $-16.205$ $-21.892$ $-46.755$ $-106.536$
% Infeasible $97\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.1316$ $0.1398$ $0.1307$ $0.3320$ $0.1697$ $0.3288$ $1.4590$
N1 Obj Mean $0.759$ $151.2e$ 15 $104.4e$ 10 $171.7e$ 17 $181.7e$ 18
Obj Median $0.455$ $-2.033$ $9.581$ $172.4e$ 17 $181.5e$ 18
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1301$ $0.1322$ $0.1316$ $0.2810$ $43.9388$
RL Obj Mean $0.493$ $-2.481$ $-4.660$ $-14.155$ $-17.195$ $-37.761$ $-90.205$
Obj Median $0.301$ $-2.756$ $-4.685$ $-14.163$ $-17.149$ $-37.761$ $-90.205$
% Infeasible $58\%$ $93\%$ $52\%$ $44\%$ $82\%$ $95\%$ $100\%$
Time (Sec) $0.0007$ $0.0007$ $0.0007$ $0.0011$ $0.0008$ $0.0009$ $0.0008$
RS Obj Mean $0.941$ $-1.097$ $-3.945$ $-11.982$ $-11.003$ $-24.598$ $-57.312$
Obj Median $0.777$ $-1.409$ $-3.956$ $-12.002$ $-11.205$ $-24.566$ $-57.312$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0014$ $0.0018$ $0.0020$ $0.0015$ $0.0020$ $0.0043$ $0.0038$
The convex quadratic problems used in the experiments are based on Donti et al. (2021), which also experimented with learning under hard (but not integer) constraints. Several modifications were introduced to ensure compatibility with our discrete version. First, all the decision variables all restricted to be integers. Second, the coefficients were rescaled. Furthermore, the original inequality constraints were removed, and the original equality constraints were transformed into inequalities:
$\min_{x\in\mathbb{Z}^{n}}\ \frac{1}{2}\bm{x}^{\intercal}\bm{Q}\bm{x}+\bm{p}^{\intercal}\bm{x}\;\text{subject to }\bm{A}\bm{x}\leq\bm{b}$
where the coefficients $\bm{Q}\in\mathbb{R}^{n\times n}$, $\bm{p}\in\mathbb{R}^{n}$, and $\bm{A}\in\mathbb{R}^{m\times n}$ were fixed and generated using random values, while $\bm{b}\in\mathbb{R}^{m}$ were treated as parametric coefficients (input features), varying across instances. Further details on the generation of $\bm{Q},\bm{p},\bm{A},$ and $\bm{b}$ can be found in Section D.1.
We experimented with quadratic problems of different sizes, from $5$ decision variables and $5$ constraints ($5\times 5$) up to $500\times 500$. The results in Table 2 summarize the performance of all methods across different problem sizes. The RC and LT methods exhibit robust performance across the board, achieving objective values second only to EX while consistently maintaining low percentages of infeasible solutions and fast solution times across all problem sizes. These methods achieve several orders of magnitude speed-ups, scaling effectively even for large instances up to $500\times 500$. Without learning-based correction, the objective values of RS are generally worse than those of RC and LT, but there are no infeasible solutions. The exact solver EX, while performing well on smaller problem sizes, fails to produce any solutions for instances of size $100\times 100$ and larger within the 60-second time limit, highlighting its limitations when handling more complex problems. N1, on the other hand, can find feasible solutions within a short time frame for smaller cases but suffers from severe numerical instability as the problem size increases. When scaled to $200\times 200$, N1 also fails to produce a solution. Both the RR and RL methods, which rely on rounding relaxations of either the solver or the prediction model, encounter significant feasibility challenges. Overall, this analysis underscores that learning-based methods like RC and LT offer considerable advantages in both solution quality and computational speed, especially for large-scale problems, compared to exact solvers or other heuristics.
It is important to note that some of the Obj Mean and Median values are extremely large. This occurs when the baseline methods, such as EX and N1, generate poor-quality feasible solutions, particularly for larger problem instances. Since the decision variables are not explicitly upper/lower bounded, the baselines occasionally produce trivial yet suboptimal solutions, leading to inflated objective values. This issue is not limited to this particular case but also appears in other problem instances, further underscoring the limitations of the baseline methods in handling larger-scale optimization tasks effectively.
In addition to evaluating solution quality, feasibility, and solving/inference times, we also measured the offline training times for our two approaches on different problem sizes. These results, along with training times for other problem types, are presented in Appendix F, where it is evident that the training times for the learning-based methods scale well with problem size.
6.2 Simple Non-convex Problem
Table 3: Results for a Simple Non-convex Problem. See the caption of Table 2 for details. “Unsolved” denotes the percentage of instances that EX or N1 could not solve within 60 seconds and is omitted for other methods, which either solve all or none.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500
RC Obj Mean $0.202$ $0.988$ $0.397$ $0.678$ $0.974$ $1.688$ $-0.716$
Obj Median $0.163$ $0.676$ $0.384$ $0.649$ $0.928$ $1.688$ $-0.717$
% Infeasible $0\%$ $2\%$ $1\%$ $2\%$ $1\%$ $0\%$ $4\%$
Time (Sec) $0.0024$ $0.0028$ $0.0037$ $0.0025$ $0.0036$ $0.0032$ $0.0061$
LT Obj Mean $0.211$ $0.605$ $-0.000$ $0.424$ $1.038$ $0.514$ $-1.646$
Obj Median $0.143$ $0.349$ $-0.018$ $0.410$ $1.014$ $0.464$ $-1.656$
% Infeasible $0\%$ $3\%$ $0\%$ $0\%$ $1\%$ $1\%$ $0\%$
Time (Sec) $0.0030$ $0.0036$ $0.0035$ $0.0033$ $0.0032$ $0.0033$ $0.0068$
EX Obj Mean $-0.027$ $-0.142$ $1.936$ $18.874$
Obj Median $-0.058$ $-0.317$ $0.288$ $11.729$
% Infeasible $0\%$ $2\%$ $2\%$ $11\%$
% Unsolved $0\%$ $3\%$ $73\%$ $77\%$
Time (Sec) $0.2046$ $4.0529$ $18.7699$ $19.5122$
RR Obj Mean $-0.047$ $-0.149$ $27.721$ $-0.782$
Obj Median $-0.089$ $-0.319$ $-0.3481$ $-0.711$
% Infeasible $63\%$ $87\%$ $98\%$ $100\%$
Time (Sec) $0.7142$ $1.5372$ $50.6019$ $54.4754$
N1 Obj Mean $1.272$ $4.651$ $29.918$ $289.992$
Obj Median $0.286$ $0.371$ $0.948$ $17.727$
% Infeasible $0\%$ $2\%$ $2\%$ $11\%$
% Unsolved $0\%$ $2\%$ $1\%$ $57\%$
Time (Sec) $0.0926$ $1.5369$ $1.8148$ $11.4497$
RL Obj Mean $0.179$ $3.282$ $5.158$ $16.626$ $19.653$ $41.756$ $93.030$
Obj Median $0.157$ $3.079$ $5.123$ $16.785$ $19.591$ $41.756$ $93.030$
% Infeasible $58\%$ $93\%$ $52\%$ $44\%$ $78\%$ $95\%$ $100\%$
Time (Sec) $0.0008$ $0.0006$ $0.0009$ $0.0010$ $0.0010$ $0.0010$ $0.0009$
RS Obj Mean $0.237$ $3.819$ $4.694$ $15.280$ $21.124$ $40.595$ $104.886$
Obj Median $0.186$ $3.503$ $4.730$ $15.156$ $20.874$ $40.639$ $104.886$
% Infeasible $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.0019$ $0.0027$ $0.0021$ $0.0026$ $0.0020$ $0.0022$ $0.0040$
To explore the performance on non-convex optimization tasks, Donti et al. (2021) extended the convex quadratic programming problem by adding a trigonometric term to the objective function. Similarly, we modified the formulation from the convex quadratic as follows:
$\min_{x\in\mathbb{Z}^{n}}\ \frac{1}{2}\bm{x}^{\intercal}\bm{Q}\bm{x}+\bm{p}^{\intercal}\sin{(\bm{x})}\;\text{subject to }\bm{A}\bm{x}\leq\bm{b}$
where the sine function is applied element-wise to the decision variables $\bm{x}$. This introduces non-convexity into the problem, making it more challenging compared to the convex case. The coefficients $\bm{Q}$, $\bm{p}$, $\bm{A}$, and $\bm{b}$ are generated in the same way as in the quadratic formulation (Section D.1). In addition, the scales of the problem and the experiment setting are also identical to those of the quadratic problems.
The results presented in Table 3 reflect patterns similar to those observed in the quadratic problem. However, the sine function exacerbates the non-convexity of the problem, rendering it more challenging for traditional methods. Despite this added complexity, the RC and LT methods perform robustly, scaling to large instances for which the baselines fail to produce any solutions.
6.3 Multi-Dimensional Mixed-Integer Rosenbrock Problem
The parametric, high-dimensional mixed-integer Rosenbrock problem serves as a challenging testbed. Adapted from the classic Rosenbrock function, this formulation introduces additional complexity by incorporating integer variables, non-linear constraints, and parametric variations in high dimensions. The problem can be expressed as follows:
$\min_{\bm{x}\in\mathbb{R}^{n},\bm{y}\in\mathbb{Z}^{n}}\quad$ $(\bm{a}-\bm{x})^{\intercal}(\bm{a}-\bm{x})+50(\bm{y}-\bm{x}^{2})^{\intercal}(\bm{y}-\bm{x}^{2})$
subject to $\|\bm{x}\|_{2}^{2}\leq nb,\mathbf{1}^{\intercal}\bm{y}\geq\frac{nb}{2},\bm{p}^{\intercal}\bm{x}\leq 0,\bm{q}^{\intercal}\bm{y}\leq 0,$
where $\bm{x}\in\mathbb{R}^{n}$ are continuous decision variables and $\bm{y}\in\mathbb{Z}^{n}$ are integer decision variables. The vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance, while the parameters $b$ and $\bm{a}$ vary. Details on how the parameters $\bm{p},\bm{q},\bm{a},$ and $b$ were generated can be found in Section D.2.
We conducted experiments on mixed-integer Rosenbrock problems with the number of decision variables ranging from 2 to 20,000; the number of constraints was fixed at 5. The results in Table 4 show that RC and LT exhibit strong performance, even outperforming EX in smaller cases, while RS performs slightly worse than RC and LT in terms of objective value. However, as the problem size increases to 10,000 variables, a noticeable decline in feasibility is observed for both RC and LT, while solver-based methods such as EX, N1, and RR fail to produce any solutions. As seen in previous experiments, RR and RL, which rely on rounding relaxations, continue to suffer from significant infeasibility issues.
Table 4: Results for the Mixed-Integer Rosenbrock Problem. The number of decision variables varies from 2 to 20,000, while the number of constraints is 5. See the caption of Table 2 for details.
Method Metric 2×5 20×5 200×5 2000×5 20000×5
RC Obj Mean $23.30$ $56.99$ $505.47$ $6245.15$ $70098.23$
Obj Median $21.22$ $50.96$ $439.79$ $6500.15$ $68253.48$
% Infeasible $2\%$ $1\%$ $1\%$ $2\%$ $32\%$
Time (Sec) $0.0031$ $0.0033$ $0.0035$ $0.0042$ $0.0856$
LT Obj Mean $23.02$ $55.06$ $608.47$ $6224.79$ $49456.67$
Obj Median $21.27$ $54.96$ $600.25$ $6097.23$ $44890.08$
% Infeasible $4\%$ $0\%$ $0\%$ $3\%$ $33\%$
Time (Sec) $0.0031$ $0.0032$ $0.0035$ $0.0047$ $0.0843$
EX Obj Mean $18.11$ $63.58$ $59.0e$ 5 $101.0e$ 10
Obj Median $18.34$ $56.84$ $28.9e$ 2 $9235.17$
% Infeasible $0\%$ $0\%$ $0\%$ $3\%$
Time (Sec) $0.4512$ $60.1385$ $59.5281$ $60.1545$
RR Obj Mean $21.52$ $52.03$ $53.3e$ 5 $68.6e$ 6
Obj Median $22.36$ $50.92$ $500.55$ $7189.05$
% Infeasible $37\%$ $42\%$ $36\%$ $55\%$
Time (Sec) $0.6756$ $1.0860$ $3.1870$ $20.8632$
N1 Obj Mean $43.86$ $88.29$ $083.0e$ 08 $132.1e$ 13
Obj Median $28.37$ $78.59$ $961.20$ $9235.17$
% Infeasible $0\%$ $0\%$ $0\%$ $3\%$
Time (Sec) $0.0752$ $0.1647$ $0.3949$ $8.7531$
RL Obj Mean $58.36$ $64.23$ $616.06$ $6186.16$ $58414.16$
Obj Median $58.02$ $62.80$ $616.91$ $5787.59$ $54241.67$
% Infeasible $86\%$ $45\%$ $55\%$ $17\%$ $26\%$
Time (Sec) $0.0006$ $0.0010$ $0.0009$ $0.0010$ $0.0007$
RS Obj Mean $25.07$ $65.38$ $634.95$ $6566.85$ $72904.02$
Obj Median $25.28$ $63.86$ $619.46$ $6441.98$ $72533.57$
% Infeasible $0\%$ $2\%$ $0\%$ $0\%$ $35\%$
Time (Sec) $0.0018$ $0.0016$ $0.0017$ $0.0027$ $0.0066$
6.4 Effect of Increasing Penalty Parameters
In this section, we evaluate the effect of progressively increasing penalty parameters on the above benchmark problems. This experiment is inspired by penalty methods commonly used in optimization, such as the Augmented Lagrangian Method (Hestenes, 1969). The weight of the penalty term is typically increased throughout the iterations. This gradual increase in penalty helps avoid numerical difficulties caused by overly strict constraints at the early stages while enhancing constraint satisfaction as the solution converges.
Table 5: Results for Increasing Penalty Parameters on Benchmark Problems. See the caption of Table 2 for details.
Problem Size Method Obj Mean Obj Median % Infeasible Time (Sec)
Convex Quadratic 100×100 RC $-10.660$ $-10.697$ $1\%$ $0.0036$
LT $-11.548$ $-11.617$ $0\%$ $0.0036$
200×200 RC $-25.350$ $-25.412$ $0\%$ $0.0033$
LT $-25.110$ $-25.297$ $0\%$ $0.0033$
500×500 RC $-60.040$ $-60.405$ $4\%$ $0.0062$
LT $-57.507$ $-57.590$ $1\%$ $0.0063$
Simple Non-Convex 100×100 RC $1.773$ $1.718$ $0\%$ $0.0038$
LT $1.932$ $1.917$ $0\%$ $0.0035$
200×200 RC $1.662$ $1.584$ $0\%$ $0.0037$
LT $2.163$ $2.060$ $0\%$ $0.0034$
[…]
Rosenbrock 200×5 RC $519.85$ $467.59$ $0\%$ $0.0034$
LT $618.51$ $624.10$ $0\%$ $0.0031$
2000×5 RC $7276.96$ $7460.26$ $0\%$ $0.0042$
LT $5780.85$ $5590.29$ $0\%$ $0.0044$
20000×5 RC $63626.35$ $58977.61$ $29\%$ $0.0044$
LT $63734.97$ $64431.54$ $39\%$ $0.0044$
In this experiment, we initialized the penalty weight at $100$, with a growth rate of $1.05$. Table 5 presents the results under this approach. In general, we observe that the method successfully improves the feasibility at the cost of objective value. It reflects the trade-off between enhancing stricter constraint satisfaction and achieving better solutions. However, for the 20000×5 Rosenbrock problem, even with progressively increasing penalties, the predictor still yields many infeasible solutions.
7 Conclusion
We have introduced a new learning-based heuristic method for MINLP. Our approach includes two novel correction layers—rounding classification and learnable threshold—that enable neural networks to generate high-quality integer solutions while preserving gradient information for training through backpropagation. These layers allow us to tackle optimization tasks with discrete variables and non-linear constraints in a way that is scalable and computationally efficient. As a self-supervised approach, our method does not require collecting optimal solutions as labels, significantly reducing the time and effort typically needed for data collection. Additionally, we proposed a strategy of progressively increasing penalties to further enhance the feasibility of the predicted solutions.
Our experiments demonstrate that our learning-based methods outperform traditional solvers and other heuristics across various problem types, including convex quadratic, non-convex, and high-dimensional mixed-integer optimization problems. Despite the increasing complexity of these tasks, our methods maintain strong performance in terms of both feasibility and solution quality, particularly in high-dimensional settings where traditional approaches often fail to produce solutions within a reasonable time due to the curse of dimensionality. To our knowledge, our work is the first to tackle learning for parametric MINLPs in full generality.
Our method allows for solving large-scale parametric MINLPs heuristically, leading to significantly better solutions and computational efficiency; however, our methods do not guarantee feasibility. Future work could focus on furthering the feasibility guarantees through alternative constraint handling or solution post-processing. For specific problem classes, it may be possible to relax only a subset of the constraints into the learning loss function while optimizing over the remaining constraints directly using existing differentiable optimization layers, e.g., from Agrawal et al. (2019). Another potential extension of our work is to redesign the neural network architecture to accommodate a varying number of instance parameters and decision variables. This can be achieved by the use of set-based, permutation-equivariant architectures including but not limited to graph neural networks (Cappart et al., 2023; Dumouchelle et al., 2024).
Appendix A Gumbel-Sigmoid Trick
The Gumbel-Sigmoid trick works by adding noise to the logits $h$, in which noise is sampled from a Gumbel distribution, defined as:
$g=-\log(-\log(U)),\quad U\sim\text{Uniform}(0,1).$
The Gumbel-Sigmoid function produces a soft approximation with random perturbation:
[…]
where $g_{1}$ and $g_{2}$ are independent samples from the Gumbel distribution, and the temperature parameter $\tau$ controls the smoothness: as $\tau\rightarrow 0$, the output becomes closer to a hard binarization. In our experiments, we set $\tau=1$ for simplicity.
Appendix B Neural Network Structure and Hyperparameters
The solution mapping $\pi_{\Theta_{1}}$ used across all learning-based methods—RC, LT, and RL—consists of five fully connected layers with ReLU activations. The rounding correction network $\varphi_{\Theta_{2}}$ for RC and LT is composed of four fully connected layers, also with ReLU activations, and incorporates Batch Normalization and Dropout with a rate of $0.2$ to prevent overfitting.
The hidden layer sizes were adjusted based on the problem size. For the convex quadratic and simple non-convex problems, the hidden layer width used in the learning-based methods was scaled accordingly, increasing from $16$, $32$, $64$ up to $1024$ for the corresponding problem sizes. Smaller problems, such as $5\times 5$, used smaller hidden layers $16$, while larger problems, such as $500\times 500$, used hidden layers with widths up to $1024$ to accommodate the complexity. Similarly, for the Rosenbrock problem, the hidden layer width was scaled based on the number of variables: a width of $4$ was used for problems with $2$ variables, $16$ for problems with $20$ variables, and up to $1024$ for problems with $10,000$ variables.
The constraint penalty weight $\lambda$ was set to $100$. All networks were trained using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $64$ over $200$ epochs. Early stopping was applied based on validation performance to ensure convergence without overfitting.
Appendix C Details for Ablation Study
C.1 Rounding after Learning
Rounding after Learning serves as an ablation study, where direct rounding is applied to the relaxed solution generated by the learning-based model $\pi_{\Theta_{1}}(\cdot)$. In this case, the loss function minimizes the objective based on the continuous relaxed solution $\bar{\bf x}\in\mathbb{R}^{n_{r}+n_{z}}$. However, when direct rounding is applied, the resulting integer solution can cause significant deviations in the objective value and may fail to satisfy feasibility constraints. This highlights the importance of end-to-end learning, where the model receives updates directly from the ultimate loss function.
C.2 Rounding with Straight-through Estimator
As another part of our ablation study, we then implement the STE to handle the non-differentiable rounding operations: 1) In this approach, the relaxed solution $\bar{\bf x}^{i}_{z}$ for integer variables is rounded down, and the STE technique enables effective gradients to pass through the rounding step during backpropagation. 2) To determine the rounding direction ${\bf b}$, we employ the Gumbel-Sigmoid function to provide approximation gradients for backward passes and output a binarized result during the forward passes.
Algorithm 2 Rounding with STE for Learning-to-optimize MINLPs: Forward Pass.
1:Training instance ${{\bm{\xi}}^{i}}$ and neural networks $\pi_{\Theta_{1}}(\cdot)$
2:Predict a continuously relaxed solution $\bar{\bf x}^{i}\leftarrow\pi_{\Theta_{1}}({{\bm{\xi}}^{i}})$
3:Round integer variables down: $\hat{{\bf x}}^{i}_{z}\leftarrow\lfloor\bar{{\bf x}}^{i}_{z}\rfloor$
[…]
5:Update integer variables: $\hat{{\bf x}}^{i}_{z}\leftarrow\hat{{\bf x}}^{i}_{z}+{\bf b}^{i}$
6:return $\hat{\bf x}^{i}$
While this method does not involve any learning mechanism for rounding correction $\varphi$, the model in this ablation setup directly rounds to the nearest integer without leveraging information from the whole relaxed solution and problem coefficients. This helps to measure how learning-based rounding can refine the rounding process and enhance overall performance.
Appendix D Detailed Problem Setup and Parameter Sampling
D.1 Convex Quadratic and Simple Non-convex Problems
For the convex and non-convex quadratic problems, $\bm{Q}$ is a diagonal matrix with entries drawn from a uniform distribution over $[0,0.01]$ to guarantee the convexity of the problem. The $\bm{p}$ was generated with entries drawn from a uniform distribution over $[0,0.1]$. The matrix $\bm{A}$ was drawn from a normal distribution with a standard deviation of $0.1$. The parameter $\bm{b}\in\mathbb{R}^{m}$, the right-hand side vector of the inequality constraints, is sampled from a uniform distribution over the range $[-1,1]$.
D.2 Ronsenbrock Problems
For the Ronsenbrock Problems, the vectors $\bm{p}\in\mathbb{R}^{n}$ and $\bm{q}\in\mathbb{R}^{n}$ are fixed for each instance and generated from a standard normal distribution.
The parameter $b$ is uniformly distributed over $[1,8]$ for each instance, and the parameter $\mathbf{a}\in\mathbb{R}^{n}$ represents a vector where elements drawn independently from a uniform distribution over $[0.5,4.5]$. The parameters $b$ and $\mathbf{a}$ influence the shape of the feasible region and the landscape of the objective function, serving as input features to the neural network.
Appendix E Experiments on Integer Linear Programs
In addition to the experiments on quadratic and non-convex problems described in the main text, we conducted further experiments on integer linear programming (ILP) problems. These experiments followed the same setup as the convex quadratic problems, except the quadratic terms were removed. Additionally, we imposed upper and lower bounds on the continuous decision variables, restricting $x$ to lie within the range $[-30,30]$:
$\min_{x\in\mathbb{Z}^{n}}\ \bm{p}^{\intercal}\bm{x}\;\text{subject to }\bm{A}\bm{x}\leq\bm{b},-30\leq\bm{x}\leq 30$
The ILP instances were generated with random coefficients for the objective function and constraint matrices in a manner consistent with the setup for the convex quadratic problems. Each instance featured varying right-hand side values, which were treated as input features. More details on the parameter generation process can be found in Appendix D.1.
Table 6: Result for an ILP Problem. Each problem size is evaluated on a test set of 100 instances. “Obj Mean” and “Obj Median” represent the mean and median objective values for this minimization problem, with smaller values being better. “Infeasible” denotes the fraction of infeasible solutions, and “Time (Sec)” is the average solving/inference time per instance. The “—” symbol indicates that no solution is found for any instance within 60 sec.
Method Metric 5×5 10×10 20×20 50×50 100×100 200×200 500×500
RC Obj Mean $0.477$ $-3.535$ $-13.335$ $-52.914$ $-71.006$ $-143.272$ $-344.111$
Obj Median $0.444$ $-3.094$ $-13.346$ $-53.119$ $-71.227$ $-143.775$ $-344.111$
% Infeasible $0\%$ $25\%$ $4\%$ $1\%$ $1\%$ $6\%$ $3\%$
Time (Sec) $0.0027$ $0.0028$ $0.0024$ $0.0034$ $0.0032$ $0.0034$ $0.0061$
LT Obj Mean $0.369$ $-1.244$ $-13.699$ $-52.786$ $-71.519$ $-150.723$ $-356.577$
Obj Median $0.354$ $-1.252$ $-13.721$ $-52.774$ $-71.520$ $-150.738$ $-356.575$
% Infeasible $4\%$ $24\%$ $0\%$ $0\%$ $0\%$ $3\%$ $7\%$
Time (Sec) $0.0029$ $0.0031$ $0.0028$ $0.0031$ $0.0034$ $0.0038$ $0.0063$
EX Obj Mean $-0.071$ $-6.543$ $-15.148$ $-56.341$ $-81.422$
Obj Median $-0.049$ $-7.096$ $-15.180$ $-56.374$ $-81.427$
% Infeasible $0\%$ $5\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1303$ $0.1689$ $0.3773$ $61.2492$ $64.1954$
RR Obj Mean $-0.145$ $-6.706$ $-15.194$ $-56.505$ $-82.608$ $-172.916$ $-422.390$
Obj Median $-0.138$ $-7.227$ $-15.226$ $-56.511$ $-82.692$ $-172.927$ $-422.397$
% Infeasible $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$ $100\%$
Time (Sec) $0.1248$ $0.1289$ $0.1307$ $0.1461$ $0.1714$ $0.4149$ $1.9499$
N1 Obj Mean $2.533$ $-2.738$ $-3.110$ $-35.219$ $7.424$
Obj Median $2.563$ $-2.526$ $-2.971$ $-35.514$ $8.883$
% Infeasible $0\%$ $5\%$ $0\%$ $0\%$ $0\%$
Time (Sec) $0.1241$ $0.1335$ $0.1362$ $0.1517$ $0.6259$
RL Obj Mean $0.158$ $-3.698$ $-14.004$ $-52.839$ $-74.392$ $-154.078$ $-379.467$
Obj Median $0.135$ $-3.429$ $-14.019$ $-52.862$ $-74.392$ $-154.061$ $-379.467$
% Infeasible $56\%$ $81\%$ $21\%$ $6\%$ $42\%$ $81\%$ $99\%$
Time (Sec) $0.0007$ $0.0008$ $0.0007$ $0.0009$ $0.0007$ $0.0009$ $0.0009$
Table 6 summarizes the results of the ILP experiments, covering problem sizes ranging from $5$ decision variables and $5$ constraints ($5\times 5$) up to $500\times 500$ with the same 60-second time limit. Overall, the learning-based methods (RC and LT) still demonstrated good performance in terms of feasibility and objective values compared to the baseline methods. In particular, the exact solver (EX, Gurobi) struggled with larger problem instances, while our methods maintained efficiency and scalability. This may be due to the random constraint matrices being dense.
However, an anomaly was observed for the learning-based methods in the $10\times 10$ case. Both RC and LT exhibited a relatively high proportion of infeasible solutions, with RC having 25% and LT having 24% infeasibility rates. This deviation stands out given the otherwise strong performance across other problem sizes, suggesting that these models may encounter difficulties when applied to some instances with specific parameter configurations. Generally, MILP solvers should be used for MILP problems. These experiments simply serve to show that our methods could be applied to MILPs.
Appendix F Training Time Comparison
In this section, we present the training times for the LR, LT, and RL methods across various problem sizes. All training runs were conducted using datasets of 9,000 instances for each problem with 1,000 instances reserved for validation per epoch. It is important to note that while the training process was set for 200 epochs, an early stopping strategy was applied, allowing the training to terminate earlier when performance plateaued.
Table 7: Training Times (in seconds) for LR, LT, and RL methods across different problem sizes for the Convex Quadratic Problem. Each method was set to train for 200 epochs, with early stopping applied.
Method 5x5 10x10 20x20 50x50 100x100 200x200 500x500
RC 242.28 225.38 153.98 237.11 141.15 149.43 606.23
LT 217.01 225.38 154.33 158.61 128.86 139.17 458.62
RL 213.53 63.96 73.72 61.95 85.91 88.49 304.80
Table 8: Training Times (in seconds) for RC, LT, and RL methods across different problem sizes for the Simple Non-convex Problem. Each method was set to train for 200 epochs, with early stopping applied.
Method 5x5 10x10 20x20 50x50 100x100 200x200 500x500
RC 257.28 144.46 173.02 138.53 136.01 104.05 116.01
LT 226.09 260.34 104.35 88.41 111.38 89.24 230.52
RL 111.07 75.67 79.28 58.86 81.43 84.28 149.87
Table 9: Training Times (in seconds) for RC, LT, and RL methods across different problem sizes for the Rosenbrock Problem. Each method was set to train for 200 epochs, with early stopping applied.
Method 2×5 20×5 200×5 2000×5 20000×5
RC 230.68 112.35 75.49 106.76 5227.05
LT 126.60 125.11 86.43 84.61 6508.41
RL 39.79 98.12 103.38 61.30 1920.59
Table 10: Training Times (in seconds) for RC, LT, and RL methods across different problem sizes for the MILP Problem. Each method was set to train for 200 epochs, with early stopping applied.
Method 5x5 10x10 20x20 50x50 100x100 200x200 500x500
RC 213.35 274.80 115.68 169.82 104.95 73.87 149.78
LT 270.57 97.47 148.06 152.03 132.28 163.37 234.96
RL 114.89 68.97 70.58 69.90 86.46 54.87 144.26
Tables 7, 8, 9 and 10, summarize the training times (in seconds) required by each method for problems of different scales. These results offer a clear view of the computational demands as problem complexity grows.